OpenStudy (cloverracer):

Need help with algebra 2 Question: Write an equation for the quadratic function (1,5)(2,8)(3,9)(4,8)

OpenStudy (ybarrap):

So a quadratic equation is an equation or order 2. This means you will have an x^2 term: $$ y=ax^2+bx+c $$ We need to find a,b and c. We will use the 3 points to generate 3 equations. With 3 equations and 3 unknowns (a,b and c) we can solve. Using the first point (1,5) we have x=1 and y=5 $$ 5=a1^2+b\cdot 1+c=a+b+c $$ Now do this for the next two points to find the other two equations. Can you do this?

OpenStudy (cloverracer):

\[8=a2^2+b \times 2+c=a+b+c\]

OpenStudy (cloverracer):

\[9=a3^2+b \times 3+ c = a+b+c\]

OpenStudy (ybarrap):

$$ 8=a2^2+b \times 2+c=4a+2b+c\\ 9=a3^2+b \times 3+ c = 9a+3b+c $$ So now you have 3 equations: $$ 5=a+b+c\\ 8=4a+2b+c\\ 9=9a+3b+c $$ You can solve for a,b and c to get your answer: $$ y=ax^2+bx+c $$ The last point (4,8) is redundant because you only need 3 points to perfectly fit a quadratic equation. Three points is sufficient because you have three unknowns, a, b and c.